# Surrogate model for Bayesian optimal experimental design in chromatography

### v1

@@ Abstract
We applied Bayesian Optimal Experimental Design (OED) in the estimation of parameters involved in the Equilibrium Dispersive Model for chromatography with two components with the Langmuir adsorption isotherm. The coefficients estimated were Henry’s coefficients, the total absorption capacity and the number of theoretical plates, while the design variables were the injection time and the initial concentration. The Bayesian OED algorithm is based on nested Monte Carlo estimation, which becomes computationally challenging due to the simulation time of the PDE involved in the dispersive model. This complication was relaxed by introducing a surrogate model based on Piecewise Sparse Linear Interpolation. Using the surrogate model instead the original reduces significantly the simulation time and it approximates the solution of the PDE with high degree of accuracy. The estimation of the parameters over strategical design points provided by OED reduces the uncertainty in the estimation of parameters. Additionally, the Bayesian OED methodology indicates no improvements when increasing the number of measurements in temporal nodes above a threshold value.
@@ 1 Introduction
Chromatography is a method to separate chemical compounds in complex mixtures by using a column packed with a selective adsorbent material. It is utilized in vastly different scales; the amounts that can be processed range from nanograms in analytical chemistry to hundreds notref, or thousands of tons per year in process industry. The separation is based on differences in propagation velocities of the compounds as they are eluted with a solvent that is pumped through the bed. Molecular level interactions with the adsorbent lead to a local distribution equilibrium (adsorption equilibrium) between the mobile fluid phase and the stationary adsorbent phase. Adsorption equilibrium is thus characteristic for the solution and the adsorbent but, at constant temperature, it depends on the solute concentrations only. The retention of the compounds in the column can be explained and predicted if the adsorption equilibrium isotherm is known.
@@
Methodologically, a series of chromatographic experiments are undertaken to empirically identify key parameters, both for the adsorption isotherms and for the non-idealities in fluid flow, termed dispersion. Predictions of a dynamic simulation model are matched with the experimental data that is presented as a time-series of concentration values obtained at the chromatographic column’s outlet. Yet, this approach is not without its challenges. Typically, a single experiment can require an hour, and real-time, high-accuracy analysis may not be possible with online detectors. Consequently, researchers may have to rely on off-line analyses that are even more time-consuming. Moreover, the errors produced by the numerical scheme used to model the chromatographic system can be significant [34]. On the other hand, simplified isotherm models with less parameters have been observed to yield more reliable (albeit physically less meaningful) parameters with the inverse method [43].
@@
Mathematical description of chromatography process is typically based on systems of hyperbolic or parabolic PDEs. In this study, we focus on the Equilibrium Dispersive Model (EDM), which comprises a nonlinear parabolic equation with an adsorption term governed by the Langmuir equations. This model depends on parameters such as the number of theoretical plates, the Langmuir constants, and Henry’s coefficient [20]. Two common approaches for the numerical solution of this problem are the finite volume methods (FVM) and the discontinuos Galerkin method (DG) [31, 32, 33]. One of the FVM methods suggested in [32] is the Koren scheme [39], which will be employed in our study.
@@
Optimal experimental design (OED) originally emerged in the context of frequentist statistics and expanded to Bayesian approach (see e.g. [16, 19, 49]). OED has an increasing role in chromatography in order to maximize the efficiency of the experiment, and to reduce waste and cost [27]. Majority of previous literature considers experimental design through empirical response models, obtained by fitting the model to measurement data, see e.g. [7, 52]. In this work, we leverage the underlying physico-chemical model of the system to produce more precise analysis of the design problem. Consequentially, our approach requires substantially larger computational effort. We note that, to our knowledge, OED utilizing a physico-chemical model for chromatography has been considered previously only in the frequentist context with D-OED criteria in [50].
Bayesian OED aims to recover design that maximizes a given utility averaged over the Bayesian joint distribution. Various different utilities have been proposed, with the most prevalent being alphabetic criteria, notably the A- and D-optimality. The A-optimality is based in the minimization of the posterior variance [1, 2, 3]. The D-optimality criteria is connected to maximizing information gain when updating prior into posterior distribution [45, 10]. In this work, we focus on D-optimality, also known as expected information gain, which for linear Gaussian problems reduces to minimizing the determinant of the posterior covariance matrix.
@@
Minimizing computational effort in Bayesian OED is a subject of active research. One avenue of research seeks to utilize a Gaussian approximation, such as the Laplace approximation, for the posterior distribution. However, common error bounds for this approach are expressed in the asymptotic limit of repetitive measurements or small noise. Additionally, the approximation requires the Jacobian of the forward mapping, which is not always available (see, e.g. [9, 15, 46, 47, 48, 25]). Another line of work aims to approximate the forward mapping or the likelihood density using a surrogate that is fast to evaluate but maintains high accuracy. It’s worth noting that the stability of such surrogate modelling was recently demonstrated in [18]. Among the various surrogate approaches, we highlight Polynomial Chaos Expansion (PCE) over sparse grids studied in [63].
@@ 1.2 Structure of the paper
The paper is organized as follow. Section 2 describes the governing equations of the EDM in chromatography, and the FVM algorithm chosen. In the same section we introduce the concept of Bayesian inverse problems, the mathematical formulation of the parameters estimation in the chromatography model, and the sampling techniques that are used for inversion. In section 3 we formalize the definition of D-OED. After that we explain how the surrogate model based in PSLI, and mathematical properties. Additionally we explain the numerical algorithms for evaluate the utility function and their order order of convergence. The section 4 we define the numerical values of the model parameters and design variables, with which we create synthetic data for different numbers of samples in the sensor. Finally the simulation is done with the true model and the surrogate model, the results are compared in both cases. In section 5 we conclude the results presented in the analysis.
@@ 2.1 Dispersive model for chromatography
In a chromatography process, dispersive models are particularly valuable for describing the behavior of components within complex mixtures as they travel through a chromatographic column. In them, the system is characterized by a set of components $c_{i}$, where $i$ ranges from 1 to $N_{c}$, representing the different species or compounds of interest. These components are governed by a set of non-linear hyperbolic partial differential equations (PDEs). Each component $c_{i}$ is described by the following governing equation
@@
where $F$ is the volumetric flow rate, $q_{i}$ models the isotherms of the system, $u$ is the linear velocity of the mobile phase and $D_{app}$ is the axial dispersion coefficient. The equation (2.1) is subjected to the zero initial conditions
@@
Boundary conditions describing the inflow and outflow of fluids are specified as follows
@@
where $\Upsilon=\frac{uT}{L}$. Then, the problem is transformed to
@@
With this change of variables, the transformed model depends on $u$ only through the new time variable $\tau$ and it simplifies the structure of the model.
@@ 2.2 Numerical implementation by the Koren scheme
The Koren scheme is a FVM method based on a piecewise-polynomial interpolation for the flux, and a flux limiter proposed by Sweby [58]. Consider regular grid $y_{m}$, $m=0,...,M$ on a unit interval with $y_{0}=0$ and $y_{M}=1$. We denote the mid-points of the grid by $y_{m+1/2}$ for $m=0,...,M-1$, and the stepsize by $\Delta y$, respectively. For what follows, let us also denote $\bar{c}=(c_{i})_{i=1}^{N_{c}}:(0,1)\times(0,\Upsilon)\to\mathbb{R}^{N_{c}}$, and the flux as $f(\bar{c})=\bar{c}.$ By chain rule, we have
$\frac{\partial q_{i}}{\partial\tau}(y,\tau)=F\sum_{j=1}^{N_{c}}\frac{dq_{i}}{dc_{j}}(y,\tau)\frac{\partial c_{j}}{\partial\tau}(y,\tau)=F(\nabla q_{i})(y,\tau)\cdot\frac{\partial\bar{c}}{\partial\tau}(y,\tau)$
@@
The Koren scheme is then given by
@@
where
$f_{\frac{1}{2}}(\tau)=\bar{c}(0,\tau),\quad f_{\frac{3}{2}}(\tau)=\bar{c}(y_{1},\tau),\qquad f_{N+\frac{1}{2}}(\tau)=\bar{c}(y_{N},\tau),$
@@
represents the ratio between consecutive fluxes with some small $\eta>0$ avoid numerical singularities. Furthermore, we utilize the central approximation
@@
The Koren scheme has spatial consistency order 2, i.e., $\|\bar{E}(\tau)\|\leq\mathcal{O}(|\Delta y|^{2})$ (see e.g. [31, 33, 39]). Following the methodology in [31], we use Runge-Kutta 4 (RK4) for temporal discretization due it is well known that RK4 has consistency order 5.
@@ 2.3 Bayesian inference
$p(\theta\mid{\bf c},d)=\frac{1}{Z({\bf c})}\exp\left(-\frac{\left\|\mathcal{G}(\theta;d)-{\bf c}\right\|^{2}}{2\sigma^{2}}\right)p_{0}(\theta),$
@@
To explore the posterior posterior distribution, Markov Chain Monte Carlo Methods (MCMC) are commonly employed, as they can generate samples from the posterior distribution without the need for the normalization constant, as discussed in [53]. Producing a sample size that accurately represents the final distribution may demand substantial computational effort. Various adaptive techniques have been developed, as discussed in [17, 22, 23]. One pragmatic and efficient approach is the DRAM algorithm, as detailed in [21], which is supported by a MATLAB toolbox and has demonstrated satisfactory results.
@@ 3 Optimal experimental design
In this section, we formulate Bayesian approach to experimental design. The uncertainty variables are $\theta=(b_{1},b_{2},Q_{s},N_{tp})$, the design variables are $d=(\tau^{inj},c^{Feed})$ and the measurements are given by a vector of size $K=2N_{s}$ with evaluations of $c_{1}$ and $c_{2}$ in the outlet $y=1$ and at times $\tau_{1},\tau_{2},\cdots,\tau_{N_{s}}$ as
@@
Here, we consider the D-OED criteria based in the Kullback–Leibler divergence (KL) between the posterior and prior measures, i.e
@@
$U(d)=\mathbb{E}D_{KL}(p(\cdot\mid{\bf c};d)\;\|\;p_{0})=\iint_{\mathbb{R}^{K}\times\Theta}p(\theta|{\bf c};d)\log\left(\dfrac{p(\theta|{\bf c};d)}{p_{0}(\theta)}\right)d\theta p({\bf c};d)d{\bf c},$ (3.1)
@@
For non-linear forward mappings that do not have sufficient linearization properties, such as our problem here, the integral in (3.1) needs to be evaluated approximatively. While Monte Carlo based algorithms are a natural choice, they tend to require high computational effort due to the inherent nested sampling task, which we discuss more below. Before considering evaluation, let us note that the computational cost is dominated by the effort needed to evaluate the forward mapping. Well-design surrogate models can reduce this cost.
@@ 3.1 Double-loop Monte Carlo integration
where $p({\bf c};d)$ refers to the marginal density of the observation, i.e., the evidence. The computational challenge in estimating (3.2) is that the integrand function intractable (excluding special cases) and cannot be directly approximated by conventional Monte Carlo methods. In literature, the traditional approach is to apply double-loop (i.e. nested) Monte Carlo integration as follows: first, an ensemble $\{(\theta^{k},{\bf c}^{k})\}_{k=1}^{M}$ is generated from the joint distribution. Second, noting the identity
@@ 3.2 Surrogated model based in Piecewise Linear Interpolation
with a surrogate mapping ${\mathcal{G}}_{N}:\Theta\times\mathcal{D}\rightarrow\mathbb{R}^{K}$. The parameter $N$ identifies the approximation rate, which here will correspond to the discretization grid. In [18], it was demonstrated that the EIG (3.1) is stable w.r.t. perturbations in the likelihood. In particular, it was shown that the induced error in EIG evaluation is bounded by the squared approximation error ${\mathcal{G}}(\theta;d)-{\mathcal{G}}_{N}(\theta;d)$ averaged over the prior.
In this work, we utilize piecewise sparse linear interpolation (PSLI) [14], which is well-suited for surrogate modelling of mappings with limited smoothness. For their rigorous construction, let us begin by introducing a one-dimensional piecewise linear interpolant $U[f]$ of a function $f:[-1,1]\rightarrow\mathbb{R}$. This interpolant is defined over a set of nodes $-1=x_{0}<x_{1}<x_{2}<\cdots<x_{m}=1$ as follows:
@@
$a_{j}(x)=\begin{cases}1-\frac{n\left|x-x_{j}\right|}{2},&\text{ if }\left|x-x_{j}\right|<\frac{2}{n}\\ 0,&\text{ otherwise. }\end{cases}$
@@
$=$ $\sum_{q-r+1\leq|\mathbf{i}|\leq r}(-1)^{r-|\mathbf{i}|}\cdot\left(\begin{array}[]{c}r-1\\ q-|\mathbf{i}|\end{array}\right)\cdot\left(U^{i_{1}}\otimes\cdots\otimes U^{i_{r}}\right),$
where ${\bf i}=(i_{1},...,i_{r})\in{\mathbb{N}}^{r}$ and $|{\bf i}|=i_{1}+...+i_{r}$. For the second identity, see e.g. [60, Lemma 1]. The resulting grid consists of $N(q,r)\leq 2^{q}\frac{(q-1)^{r-1}}{(r-1)!}$ points (see e.g. [51]); the arrangement is illustrated on a two-dimensional domain in Figure 1.
@@
$F_{d}^{k}=\begin{cases}C^{k}([-1,1])&\text{ if }r=1\\ \left\{f:[-1,1]^{r}\rightarrow\mathbb{R}\mid D^{\alpha}f\text{ continuous if }\alpha_{i}\leq k\text{ for all }i\right\}&\text{ if }r>1\end{cases}$
@@
$\left\|f-S_{q,r}f\right\|_{\infty}\leq c_{r,k}N^{-k}(\log N)^{(k+1)(r-1)}\left\|f\right\|_{C^{k}},$
@@
We note that piecewise linear interpolation is not order optimal for $k>2$ and, therefore, the rate in Theorem 3 does not extend beyond $k=2$, see [8, Remark 10]. An efficient algorithmic implementation is discussed in [35, 36], and a free toolbox is available [37, 38]. Moreover, Theorem 3 does not guarantee convergence for continuous functions. Nonetheless, continuity is typically the minimum requirement for interpolation, and numerical experiments in [35, 51] show convergence for continuous benchmark functions. Finally, we define our surrogate model by constructing
@@
with ${\mathcal{G}}=({\mathcal{G}}_{k})_{k=1}^{K}$, on a grid of points $d_{j}\in{\mathcal{D}}$, where $N=N(q,r)$. The full mapping for all $d\in{\mathcal{D}}$ is obtained by linear interpolation. We note that while theoretically establishing smoothness of the forward mapping ${\mathcal{G}}$ is outside the scope of this work, no issues were encountered in our numerical simulations.

### v2

@@ Abstract
We applied Bayesian Optimal Experimental Design (OED) in the estimation of parameters involved in the Equilibrium Dispersive Model for chromatography with two components with the Langmuir adsorption isotherm. The coefficients estimated were Henry’s coefficients, the total absorption capacity and the number of theoretical plates, while the design variables were the injection time and the initial concentration. The Bayesian OED algorithm is based on nested Monte Carlo estimation, which becomes computationally challenging due to the simulation time of the PDE involved in the dispersive model. This complication was relaxed by introducing a surrogate model based on Piecewise Sparse Linear Interpolation. Using the surrogate model instead of the original one reduces significantly the simulation time and approximates the solution of the PDE with high degree of accuracy. The estimation of the parameters over strategical design points provided by OED reduces the uncertainty in the estimation of parameters. Additionally, the Bayesian OED methodology shows that there is no improvement of results after certain threshold values of injection time, concentrations and number of observation instances.
@@ 1 Introduction
Chromatography is a method to separate chemical compounds in complex mixtures by using a column packed with a selective adsorbent material. It is utilized in vastly different scales; the amounts that can be processed range from nanograms in analytical chemistry to hundreds or thousands of tons per year in process industry. The separation is based on differences in propagation velocities of the compounds as they are eluted with a solvent that is pumped through the bed. Molecular level interactions with the adsorbent lead to a local distribution equilibrium (adsorption equilibrium) between the mobile fluid phase and the stationary adsorbent phase. Adsorption equilibrium is thus characteristic for the solution and the adsorbent but, at constant temperature, it depends on the solute concentrations only. The retention of the compounds in the column can be explained and predicted if the adsorption equilibrium isotherm is known.
@@
Methodologically, a series of chromatographic experiments are undertaken to empirically identify key parameters, both for the adsorption isotherms and for the non-idealities in fluid flow, termed dispersion. Predictions of a dynamic simulation model are matched with the experimental data that is presented as a time series of concentration values obtained at the chromatographic column’s outlet. Yet, this approach is not without its challenges. Typically, a single experiment can require an hour, and real-time, high-accuracy analysis may not be possible with online detectors. Consequently, researchers may have to rely on off-line analyses that are even more time-consuming. Moreover, the errors produced by the numerical scheme used to model the chromatographic system can be significant [34]. On the other hand, simplified isotherm models with less parameters have been observed to yield more reliable (albeit physically less meaningful) parameters with the inverse method [43].
@@
Mathematical description of chromatography process is typically based on systems of hyperbolic or parabolic PDEs. In this study, we focus on the Equilibrium Dispersive Model (EDM), which comprises a nonlinear parabolic equation with an adsorption term governed by the Langmuir equations. This model depends on parameters such as the number of theoretical plates, the Langmuir constants, and Henry’s coefficient [20]. Two common approaches for the numerical solution of this problem are the finite volume methods (FVM) and the discontinuous Galerkin method (DG) [31, 32, 33]. One of the FVM methods suggested in [32] is the Koren scheme [39], which will be employed in our study.
@@
Optimal experimental design (OED) originally emerged in the context of frequentist statistics and expanded to Bayesian approach (see e.g. [16, 19, 49]). OED has an increasing role in chromatography in order to maximize the efficiency of the experiment, and to reduce waste and cost [27]. Majority of the previous literature considers experimental design through empirical response models, obtained by fitting the model to measurement data, see e.g. [7, 53]. In this work, we leverage the underlying physico-chemical model of the system to produce more precise analysis of the design problem. Consequentially, our approach requires substantially larger computational effort. We note that, to our knowledge, OED utilizing a physico-chemical model for chromatography has been considered previously only in the frequentist context with D-OED criteria in [50].
Bayesian OED aims to recover a design that maximizes a given utility averaged over the Bayesian joint distribution. Various different utilities have been proposed, with the most prevalent being alphabetic criteria, notably the A- and D-optimality. The A-optimality is based on the minimization of the posterior variance [1, 2, 3]. The D-optimality criteria is connected to maximizing information gain when updating prior into posterior distribution [45, 10]. In this work, we focus on D-optimality, also known as expected information gain, which for linear Gaussian problems reduces to minimizing the determinant of the posterior covariance matrix.
@@
Minimizing computational effort in Bayesian OED is a subject of active research. One avenue of research seeks to utilize a Gaussian approximation, such as the Laplace approximation, for the posterior distribution. However, common error bounds for this approach are expressed in the asymptotic limit of repetitive measurements or small noise. Additionally, the approximation requires the Jacobian of the forward mapping, which is not always available (see, e.g. [9, 15, 46, 47, 48, 25]). Another line of work aims to approximate the forward mapping or the likelihood density using a surrogate that is fast to evaluate but maintains high accuracy. It is worth noting that the stability of such surrogate modelling was recently demonstrated in [18]. Among the various surrogate approaches, we highlight Polynomial Chaos Expansion (PCE) over sparse grids studied in [65] and applied in Bayesian OED in [28, 29, 30, 52]. Finally we can mention other approach based in a fast calculation using sparse stochastic collocation [55].
@@ 1.2 Structure of the paper
The paper is organized as follows. Section 2 describes the governing equations of the EDM in chromatography, and the FVM algorithm chosen. In the same section we introduce the concept of Bayesian inverse problems, the mathematical formulation of the parameters estimation in the chromatography model, and the sampling techniques that are used for inversion. In Section 3 we formalize the definition of D-OED. After that we explain the surrogate model based on PSLI, and mathematical properties. Additionally we explain the numerical algorithms for evaluating the utility function and their order order of convergence. In Section 4 we define the numerical values of the model parameters and design variables, with which we create synthetic data for different numbers of samples in the sensor. Finally the simulation is done with the true model and the surrogate model, and the results are compared in both cases. In Section 5 we conclude the results presented in the analysis.
@@ 2.1 Dispersive model for chromatography
In a chromatography process, dispersive models are particularly valuable for describing the behavior of components within complex mixtures as they travel through a chromatographic column. In them, the system is characterized by a set of components $c_{i}$, where $i$ ranges from 1 to $N_{c}$, representing the different species or compounds of interest. These components are governed by a set of nonlinear hyperbolic partial differential equations (PDEs). Each component $c_{i}$ is described by the following governing equation
@@
where $F$ is the ratio of volumetric fractions of stationary and mobile phases, $q_{i}$ models the isotherms of the system, $u$ is the linear velocity of the mobile phase, $D_{app}$ is the axial dispersion coefficient, $L$ is the length of the column and $T$ is the total time simulation. The equation (2.1) is subject to the zero initial conditions
@@
The inflow and outflow of fluids in the boundaries are specified by Danckwerts conditions [20].
@@
where $\Upsilon=\frac{uT}{L}$ is a dimensionless time. Then, the problem is transformed to
@@
With this change of variables, the transformed model depends on $u$ only through the new time variable $\tau$ which simplifies the structure of the model.
@@ 2.2 Numerical implementation by the Koren scheme
The Koren scheme is a FVM method based on a piecewise-polynomial interpolation for the flux, and a flux limiter proposed by Sweby [60]. Consider regular grid $y_{m}$, $m=0,...,N_{t}+1$ on a unit interval with $y_{0}=0$ and $y_{N_{t}+1}=1$. We denote the mid-points of the grid by $y_{m+1/2}$ for $m=0,...,N_{t}$, and the stepsize by $\Delta y$, respectively.
For what follows, let us also denote $\bar{c}=(c_{i})_{i=1}^{N_{c}}:(0,1)\times(0,\Upsilon)\to\mathbb{R}^{N_{c}}$, and the flux as $f(\bar{c})=\bar{c}.$ By chain rule, we have
$\frac{\partial q_{i}}{\partial\tau}(y,\tau)=\sum_{j=1}^{N_{c}}\frac{dq_{i}}{dc_{j}}(y,\tau)\frac{\partial c_{j}}{\partial\tau}(y,\tau)=(\nabla q_{i})(y,\tau)\cdot\frac{\partial\bar{c}}{\partial\tau}(y,\tau)$
@@
For $m=1,2,\cdots,N_{t}$, the Koren scheme is given by
@@
represents the ratio between consecutive fluxes with some small $\eta>0$ to avoid numerical singularities. In the cases $m=1$ and $N_{t}$, the fluxes are calculated as follows
$f_{\frac{1}{2}}(\tau)=\bar{c}(0,\tau),\quad f_{\frac{3}{2}}(\tau)=\bar{c}(y_{1},\tau),\quad f_{N_{t}+\frac{1}{2}}(\tau)=\bar{c}(y_{N_{t}},\tau).$
Furthermore, for approximate the partial derivatives we utilize the central scheme
@@
The Koren scheme has spatial consistency order 2, i.e., $\|\bar{E}(\tau)\|\leq\mathcal{O}(|\Delta y|^{2})$ (see e.g. [31, 33, 39]). Following the methodology in [31], we use Runge-Kutta 4 (RK4) for temporal discretization since it is well known that RK4 has consistency order 5.
@@ 2.3 Bayesian inference
To explore the posterior distribution, Markov Chain Monte Carlo (MCMC) methods are commonly employed, as they can generate samples from the posterior distribution without the need for the normalization constant, as discussed in [54]. Producing a sample size that accurately represents the final distribution may demand substantial computational effort. Various adaptive techniques have been developed, as discussed in [17, 22, 23]. One pragmatic and efficient approach is the DRAM algorithm, as detailed in [21], which is supported by a MATLAB toolbox and has demonstrated satisfactory results.
@@ 3 Optimal experimental design
In this section, we formulate Bayesian approach to experimental design. Here we assume that we have only two components and the initial injected concentration is equal for both, i.e., $c_{1}^{Feed}=c_{2}^{Feed}=c^{Feed}$. Then, the uncertainty variables are $\theta=(b_{1},b_{2},Q_{s},N_{tp})$, the design variables are $d=(\tau^{inj},c^{Feed})$ and the measurements are given by a vector of size $K=2N_{s}$ with evaluations of $c_{1}$ and $c_{2}$ in the outlet $y=1$ and at times $\tau_{1},\tau_{2},\cdots,\tau_{N_{s}}$ as
@@
Here, we consider the D-OED criteria based on the Kullback–Leibler divergence (KL) between the posterior and prior measures, i.e
@@
$U(d)=\mathbb{E}^{{\bf c}}D_{KL}(p(\cdot\mid{\bf c};d)\;\|\;p_{0})=\iint_{\mathbb{R}^{K}\times\Theta}p(\theta|{\bf c};d)\log\left(\dfrac{p(\theta|{\bf c};d)}{p_{0}(\theta)}\right)d\theta p({\bf c};d)d{\bf c},$ (3.1)
@@
For nonlinear forward mappings that do not have sufficient linearization properties, such as our problem here, the integral in (3.1) needs to be evaluated approximatively. While Monte Carlo-based algorithms are a natural choice, they tend to require high computational effort due to the inherent nested sampling task, which we discuss more below. Before considering evaluation, let us note that the computational cost is dominated by the effort needed to evaluate the forward mapping. Well-designed surrogate models can reduce this cost.
@@ 3.1 Double-loop Monte Carlo integration
where $p({\bf c};d)$ refers to the marginal density of the observation, i.e., the evidence. The computational challenge in estimating (3.2) is that the integrand function is intractable (excluding special cases) and cannot be directly approximated by conventional Monte Carlo methods. In literature, the traditional approach is to apply double-loop (i.e. nested) Monte Carlo integration as follows: first, an ensemble $\{(\theta^{k},{\bf c}^{k})\}_{k=1}^{M}$ is generated from the joint distribution. Second, noting the identity
@@ 3.2 Surrogated model based in Piecewise Linear Interpolation
with a surrogate mapping ${\mathcal{G}}_{N}:\Theta\times\mathcal{D}\rightarrow\mathbb{R}^{K}$. The parameter $N$ identifies the approximation rate, which here will correspond to the discretization grid. In [18], it was demonstrated that the $U(d)$ is stable w.r.t. perturbations in the likelihood. In particular, it was shown that the induced error in EIG evaluation is bounded by the squared approximation error ${\mathcal{G}}(\theta;d)-{\mathcal{G}}_{N}(\theta;d)$ averaged over the prior.
In this work, we utilize piecewise sparse linear interpolation (PSLI) [14], which is well-suited for surrogate modelling of mappings with limited smoothness. For their rigorous construction, let us begin by introducing a one-dimensional piecewise linear interpolant $U[f]$ of a function $f:[-1,1]\rightarrow\mathbb{R}$. This interpolant is defined over a set of nodes $-1=x_{0}<x_{1}<x_{2}<\cdots<x_{m}=1$ and a nodal basis $\{a_{j}\}_{j=1}^{m}$ as follows:
@@
$a_{j}(x)=\begin{cases}1-\frac{n\left|x-x_{j}\right|}{2},&\text{ if }\left|x-x_{j}\right|<\frac{2}{m}\\ 0,&\text{ otherwise. }\end{cases}$
@@
$=$ $\sum_{q-r+1\leq|\mathbf{i}|\leq q}(-1)^{r-|\mathbf{i}|}\cdot\left(\begin{array}[]{c}r-1\\ q-|\mathbf{i}|\end{array}\right)\cdot\left(U^{i_{1}}\otimes\cdots\otimes U^{i_{r}}\right),$
where ${\bf i}=(i_{1},...,i_{r})\in{\mathbb{N}}^{r}$ and $|{\bf i}|=i_{1}+...+i_{r}$. For the second identity, see e.g. [62, Lemma 1]. The resulting grid consists of $N(q,l)\leq 2^{q}\binom{q-1}{l-1}$ points (see e.g. [51]); the arrangement is illustrated on a two-dimensional domain in Figure 1.
@@
$F_{r}^{k}=\begin{cases}C^{k}([-1,1])&\text{ if }r=1\\ \left\{f:[-1,1]^{r}\rightarrow\mathbb{R}\mid D^{\alpha}f\text{ continuous if }\alpha_{i}\leq k\text{ for all }i\right\}&\text{ if }r>1\end{cases}$
@@
$\|f-S_{q,r}f\|_{\infty}\leq c_{r,k}N^{-k}(\log N)^{(k+1)(r-1)}\|f\|_{C^{k}},$
@@
We note that piecewise linear interpolation is not order optimal for $k>2$ and, therefore, the rate in Theorem 3 does not extend beyond $k=2$, see [8, Remark 10]. An efficient algorithmic implementation is discussed in [35, 36], and a free toolbox is available [37, 38]. Moreover, Theorem 3 does not guarantee convergence for continuous functions that are not continuously differentiable. Nonetheless, continuity is typically the minimum requirement for interpolation, and numerical experiments in [35, 51] show convergence for continuous benchmark functions. Finally, we define our surrogate model over the parameter space $\Theta$ by constructing
@@
with ${\mathcal{G}}=({\mathcal{G}}_{k})_{k=1}^{K}$, on a grid of points $d_{j}\in{\mathcal{D}}$, where $N=N(q,r)$. The full mapping for all $d\in{\mathcal{D}}$ is obtained by linear interpolation. We note that, while theoretically establishing smoothness of the forward mapping ${\mathcal{G}}$ is outside the scope of this work, no issues were encountered in our numerical simulations.